### v1

@@ 4.4 ML models and interpretation
Diamond is designed to be compatible with a wide range of ML models. To demonstrate the broad applicability of Diamond, we use it in conjunction with machine learning models from methodologically distinct categories, including DNN models, tree-based models, and factorization-based models.
@@
For DNNs, we configured a pyramid-shaped multi-layer perceptron (MLP) model with the exponential linear unit activation and four hidden layers, each having neuron sizes of 2p, p, p/2, and p/4 respectively, where p denotes the input dimensionality. The model follows the guidance of [33] and includes a plugin pairwise-coupling layer, connecting each original feature with its knockoff counterpart in a pairwise fashion to maximize statistical power (See details in Sec. S.1). Using an MLP as the ML model, we employed three representative interaction importance measures to clarify the relationship between features and responses. We began with a model-specific interaction importance measure derived from the model weights, specifically designed for the MLP architecture (See details in Sec. S.2). To demonstrate Diamond’s flexibility, we also used Integrated Hessian [22] and Expected Hessian [14], two state-of-the-art, model-agnostic interaction importance measures, to elucidate the relationships between features and responses without making assumptions about any specific model architecture (See details in Sec. S.3). Both model-agnostic interaction importance measures are calculated using the Path-Explain library [22].
@@
For tree-based models, we employed two widely used gradient-boosted decision tree representatives—XGBoost [8] and LightGBM [24]. We used the implementations provided by the two libraries with default settings. Using either XGBoost or LightGBM as the ML model, we utilized TreeSHAP [37], a Shapley value-based interpretation method tailored for tree-based ML models, to elucidate the relationships between features and responses. We used the implementation provided by the SHAP library [36] to calculate both univariate and pairwise feature importance.
For factorization-based models, we used a widely adopted representative method—Factorization machines (FM) [41]. We used 2-Way FM implemented in the xLearn library [39] with default settings, which models up to second-order feature interactions. Since 2-Way FM is explicitly modelled as the weighted sum of univariate and pairwise feature interactions, we used the learned coefficients of these interactions as the corresponding importance measures.
@@ 4.5 Knockoff generation
Diamond relies on knockoffs to control the FDR. It is worth mentioning that conventional knockoffs are limited to Gaussian settings, which may not be applicable in many practical scenarios. In this study, we focus on the state-of-the-art knockoff design without assuming any specific feature distribution. In particular, we consider commonly-used non-Gaussian knockoff generation methods such as KnockoffsDiagnostics [5], KnockoffGAN [23], Deep knockoffs [42], and VAE knockoffs [31].
@@ 4.6 Alternative FDR estimation methods
We evaluate the performance of Diamond in comparison to other baseline methods. For the first baseline method, we employ a permutation-based approach to calculate the interaction-wise FDR. Specifically, this involves using a previously described permutation procedure tailored for neural networks to assess the significance of interactions and calculate permutation p-values [12], followed by the Benjamini–Hochberg procedure [4] to estimate the FDR.
@@ 4.7 Code and data availability
Appendix S.1 Knockoff-tailored multi-layer perceptron model
Figure S.1: Illustration of knockoff-tailored MLP model.
By following the guidance of [33], Diamond utilizes a knockoff-tailored multi-layer perceptron (MLP) model that includes a plugin pairwise-coupling layer, connecting each original feature with its knockoff counterpart in a pairwise fashion to maximize statistical power. Specifically, the model takes as input an augmented data matrix $(\mathbf{X},\tilde{\mathbf{X}})\in\mathbb{R}^{n\times 2p}$, created by concatenating the original data matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ with its knockoffs $\tilde{\mathbf{X}}\in\mathbb{R}^{n\times p}$. The augmented data matrix is fed to an off-the-shelf MLP model through a plugin pairwise-coupling layer composed of $p$ filters, encapsulated by $\mathbf{F}=(F_{1},\cdots,F_{p})\in\mathbb{R}^{p}$, where each $j$-th filter connects feature $X_{j}$ and its knockoff counterpart ${\tilde{X}}_{j}$, as shown in Fig. S.1.
@@
Appendix S.2 Model-specific interaction importance measure
The model-specific interaction importance measure is based on the knockoff-tailored MLP model design and notation described in Sec. S.1, Given the pairwise-coupling layer weights: $\mathbf{Z}\in\mathbb{R}^{p}$ and $\tilde{\mathbf{Z}}\in\mathbb{R}^{p}$ as well as the MLP weights in the $l$-th layer: $\mathbf{W}^{(l)}$, the model-specific interaction importance measure decomposes into two factors: (1) the relative importance between the original feature and its knockoff counterpart, encoded by concatenated filter weights $\mathbf{Z}^{\text{Agg}}=(\mathbf{Z},\tilde{\mathbf{Z}})\in\mathbb{R}^{2p}$, and (2) the relative importance among all $p$ features, encoded by the weight matrix $\mathbf{W}^{(0)}\in\mathbb{R}^{p\times p_{1}}$ and the aggregated weights $\mathbf{W}^{\text{Agg}}=\prod_{i=1}^{L}\mathbf{W}^{(i)}\in\mathbb{R}^{p_{1}}$. (See [18] for theoretical insights regarding $\mathbf{W}^{\text{Agg}}$.)
@@
Appendix S.3 Model-agnostic interaction importance measure
@@
Appendix S.4 Simulated dataset
Following [52], we use simulation datasets to evaluate Diamond. 10 simulation functions (described in Tab. S.1) are used to generate the simulation datasets.
Table S.1: A test suite of data-generating simulation functions proposed by [52].
@@
We discovered that the proposed calibration procedure; without it, the FDR cannot be controlled by naively using reported interaction importance from existing methods (Fig. S.2). To gain insight into the FDR control failure, we conducted a qualitative comparison assessing interaction importance before and after calibration using the simulation function $F_{1}$ across different ML models (Fig. S.3). The primary cause of the FDR control failure lies in the distribution disparity between original-only interactions and those involving knockoffs. This suggests a violation of the knockoff filter’s assumption in controlling FDR. The proposed calibration procedure mitigates the disparity by extracting non-additive interactive effects from the reported interaction importance measures, thereby enhancing the utility of knockoff-involving interactions as a negative control for FDR estimation.
Figure S.2: Evaluation of Diamond without calibration on simulated datasets. Diamond without the proposed calibration procedure fails to effectively control the FDR across most ML models. The only exception is Factorization machines, which explicitly model univariate and pairwise feature interactions. The interaction importance measure, represented by the learned coefficients of these interactions, should not include marginal effects to be removed, as the proposed calibration procedure does.
@@
Figure S.3: The calibration procedure mitigates the disparity between original-only interactions and those involving knockoffs. (a) The reported interaction importance from existing methods reveals a clear disparity in the distribution of interaction importance between original-only interactions and those involving knockoffs, on simulation function $F_{1}$. The proposed calibration procedure mitigates the disparity by aligning the two distributions more closely. (b) The reported interaction importance from existing methods shows a clear disparity in the cumulative density function between original-only interactions and those involving knockoffs. The proposed calibration procedure mitigates the disparity by aligning the cumulative density functions to be more similar to each other.
Figure S.4: Evaluating the feature contribution to enhancer status prediction in Drosophila enhancer dataset. Each feature contributes differently to predicting enhancer status, as measured by the Expected Gradient scores in the MLP model.
@@
Figure S.5: Evaluating Diamond on a real Drosophila enhancer dataset. (a) Diamond is applied with LightGBM model to identify important non-additive interactions. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top five interaction annotated. The annotated transcription factors are labeled by their UniProt identifiers. The red stars indicate well-characterized interactions in early Drosophila embryos as ground truth. (b) The top interactions reported by the MLP model are qualitatively evaluated from three aspects: the contribution of feature values to the marginal and interaction importance measure, and the contribution of the marginal importance measures to the interaction importance measure.
@@
Figure S.6: Evaluating Diamond on a real Drosophila enhancer dataset. (a) Diamond is applied with XGBoost model to identify important non-additive interactions. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top five interaction annotated. The annotated transcription factors are labeled by their UniProt identifiers. The red stars indicate well-characterized interactions in early Drosophila embryos as ground truth. (b) The top interactions reported by the MLP model are qualitatively evaluated from three aspects: the contribution of feature values to the marginal and interaction importance measure, and the contribution of the marginal importance measures to the interaction importance measure.

### v2

@@ 4.4 ML models
Diamond is designed to be compatible with a wide range of ML models. To demonstrate the broad applicability of Diamond, we use it in conjunction with machine learning models from methodologically distinct categories, including DNN models, Kolmogorov-Arnold networks (KANs) [36], tree-based models, and factorization-based models.
For DNNs, we selected three representative models: multi-layer perceptron (MLP), convolutional neural network (CNN), and transformer models. It is worth noting that CNNs are not well-suited for modeling tabular data and are included here as an example of model misspecification. We configured a pyramid-shaped MLP model with the exponential linear unit activation and four hidden layers, each having neuron sizes of 2p, p, p/2, and p/4 respectively, where p denotes the input dimensionality. We configured a CNN model consisting of two convolutional layers. The first layer has one input channel and 32 output channels, while the second layer has 32 input channels and 64 output channels, followed by max pooling. Batch normalization and ReLU activation are applied after each convolution. The output is then flattened and passed through two fully connected layers, mapping to 128 units before reaching the output dimension. For transformer models, we used FT-Transformer [22], a model specifically designed for tabular data. By default, it consists of 6 layers, 8 attention heads, a 32-dimensional feedforward network, and a 0.1 dropout rate. All DNN models follow the guidance of [38] and include a plugin pairwise-coupling layer, connecting each original feature with its knockoff counterpart in a pairwise fashion to maximize statistical power (See details in Sec. A.1).
For KANs, we configured a two-layer model, with a hidden layer of width p/2, followed by a mapping to the output dimension. For tree-based models, we employed two widely used gradient-boosted decision tree representatives—XGBoost [10] and LightGBM [28]. We used the implementations provided by the two libraries with default settings. For factorization-based models, we used a widely adopted representative method—Factorization machines (FM) [46]. We used 2-Way FM implemented in the xLearn library [44] with default settings, which models up to second-order feature interactions.
4.5 Model interpretation
@@
Diamond utilizes representative importance measures to interpret the trained ML models and elucidate the relationship between features and responses. For DNN and KAN models, we employed a representative model-agnostic interaction importance measure, Expected Hessian [16], to clarify the relationships between features and responses without making assumptions about any specific model architecture (See details in Sec. A.3). To demonstrate Diamond’s flexibility, we also used Integrated Hessian [25], another state-of-the-art, model-agnostic interaction importance measure, alongside a model-specific interaction importance measure derived from the model weights, specifically designed for the MLP architecture (See details in Sec. A.2). Both model-agnostic interaction importance measures, Expected Hessian and Integrated Hessian, are computed using the Path-Explain library [25].
@@
For tree-based models such as XGBoost and LightGBM, we used TreeSHAP [42], a Shapley value-based interpretation method specifically designed for tree-based ML models, to clarify the relationships between features and responses. We used the implementation provided by the SHAP library [41] to calculate both univariate and pairwise feature importance. For factorization-based models, since 2-Way FM is explicitly modeled as the weighted sum of univariate and pairwise feature interactions, we used the learned coefficients of these interactions as the corresponding importance measures.
@@ 4.6 Knockoff generation
Diamond relies on knockoffs to control the FDR. It is worth mentioning that conventional knockoffs are limited to Gaussian settings, which may not be applicable in many practical scenarios. In this study, we focused on the state-of-the-art knockoff design without assuming any specific feature distribution. In particular, we considered commonly-used non-Gaussian knockoff generation methods such as KnockoffsDiagnostics [7], KnockoffGAN [26], Deep knockoffs [48], and VAE knockoffs [35]. Furthermore, to assess Diamond’s robustness in the presence of poorly generated knockoffs, we created invalid knockoffs by independently permuting each feature across samples, thus violating the definition of knockoffs outlined in Sec. 4.1.
@@ 4.7 Alternative FDR estimation methods
We evaluate the performance of Diamond in comparison to three baseline methods. For the first two baseline methods, we employ a permutation-based approach to calculate the interaction-wise FDR. Specifically, this involves using a previously described permutation procedure tailored for neural networks to assess the significance of interactions and calculate permutation p-values [14]. The permutation p-values are combined with the Benjamini–Hochberg procedure [5] or the Benjamini-Yekutieli procedure [6] to estimate the FDR. The difference between the Benjamini–Hochberg procedure and the Benjamini-Yekutieli procedure is that the former assumes the p-values are independent or positively dependent, while the latter imposes no assumptions on the dependencies among p-values.
@@ 4.8 Code and data availability
Appendix A Appendix
Appendix A.1 Knockoff-tailored multi-layer perceptron model
Figure A.1: Illustration of knockoff-tailored MLP model.
By following the guidance of [38], Diamond utilizes a knockoff-tailored multi-layer perceptron (MLP) model that includes a plugin pairwise-coupling layer, connecting each original feature with its knockoff counterpart in a pairwise fashion to maximize statistical power. Specifically, the model takes as input an augmented data matrix $(\mathbf{X},\tilde{\mathbf{X}})\in\mathbb{R}^{n\times 2p}$, created by concatenating the original data matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ with its knockoffs $\tilde{\mathbf{X}}\in\mathbb{R}^{n\times p}$. The augmented data matrix is fed to an off-the-shelf MLP model through a plugin pairwise-coupling layer composed of $p$ filters, encapsulated by $\mathbf{F}=(F_{1},\cdots,F_{p})\in\mathbb{R}^{p}$, where each $j$-th filter connects feature $X_{j}$ and its knockoff counterpart ${\tilde{X}}_{j}$, as shown in Fig. A.1.
@@ Appendix A.1 Knockoff-tailored multi-layer perceptron model
Appendix A.2 Model-specific interaction importance measure
The model-specific interaction importance measure is based on the knockoff-tailored MLP model design and notation described in Sec. A.1, Given the pairwise-coupling layer weights: $\mathbf{Z}\in\mathbb{R}^{p}$ and $\tilde{\mathbf{Z}}\in\mathbb{R}^{p}$ as well as the MLP weights in the $l$-th layer: $\mathbf{W}^{(l)}$, the model-specific interaction importance measure decomposes into two factors: (1) the relative importance between the original feature and its knockoff counterpart, encoded by concatenated filter weights $\mathbf{Z}^{\text{Agg}}=(\mathbf{Z},\tilde{\mathbf{Z}})\in\mathbb{R}^{2p}$, and (2) the relative importance among all $p$ features, encoded by the weight matrix $\mathbf{W}^{(0)}\in\mathbb{R}^{p\times p_{1}}$ and the aggregated weights $\mathbf{W}^{\text{Agg}}=\prod_{i=1}^{L}\mathbf{W}^{(i)}\in\mathbb{R}^{p_{1}}$. (See [20] for theoretical insights regarding $\mathbf{W}^{\text{Agg}}$.)
@@ Appendix A.2 Model-specific interaction importance measure
Appendix A.3 Model-agnostic interaction importance measure
@@
Appendix A.4 FDR estimation
We estimate the FDR for non-additive interactions by following the strategy proposed by [62], which was originally designed to estimate the FDR for cross-linked peptide detection. Despite the conceptually different application backgrounds, the FDR estimation for cross-linked peptide detection and non-additive interaction detection are essentially analogous. Specifically, in the context of cross-linked peptide detection, there is no ground truth for false positive cross-linked peptides. Therefore, researchers developed methods to determine the FDR of detected cross-linked peptides in a data-driven way, even when the ground truth is not known. Researchers relied on a target-decoy strategy, where verifiably incorrect decoy sequences are appended to the target-sequence database used by the search engine. The rate of false positive hits mapping to the target database is then estimated based on the number of hits mapping to the decoy database. Here, the decoys are conceptually analogous to the knockoffs. By examining the cross-link detections, four different, equally likely cases can be observed: T-T, T-D, D-T and D-D, where T denotes a target hit and D denotes a decoy hit. The key idea proposed by [62] is to use decoy cross-links (i.e., T-D, D-T and D-D) to estimate the FDR.
Now we adopt the strategy proposed by [62] to our setting. They calculated the expected number of false positive interactions by separately estimating the counts of different types of false interactions. Specifically, three (not necessarily mutually exclusive) types of false interactions are distinguished below. Type TC–!TC interactions involve one correct target feature and one feature that is not a correct target feature. Type !TC–!TC interactions consist of two features, neither of which are correct target features. Type K–K interactions consist of two features that are both knockoff features (this type can be directly counted). Note that type K–K interactions are a subset of type !TC–!TC interactions, as knockoff features are not correct target features.
Given a set of detected interactions exceeding a certain score threshold, the total number of false positives #$\#(\text{FP})$ in this set can be decomposed as follows:
#$\#(\text{FP})=\#(\text{FP}_{\text{TC--!TC}})+\#(\text{FP}_{\text{!TC--!TC}})$ (13)
The expected count of each false positive type given the knockoff counts can be estimated as follows:
$\widehat{E}\left[\#(\text{FP}_{\text{TC--!TC}})\right]$ =#$=\#(\text{Knockoff}_{\text{TC--!TC}})\times r_{\text{TC--!TC}}$ (14)
$\widehat{E}\left[\#(\text{FP}_{\text{!TC--!TC}})\right]$ =#$=\#(\text{Knockoff}_{\text{!TC--!TC}})\times r_{\text{!TC--!TC}}$
where #$\#(\text{Knockoff}_{y})$ denotes the number of knockoff-involving interactions of type $y$, and $r_{y}$ represents the respective original-knockoff frequencies. For TC–!TC and !TC–!TC interactions we have original-knockoff frequencies $r_{\text{TC--!TC}}$ of $1:1$ and $r_{\text{!TC--!TC}}$ of $1:3$. The original-knockoff frequencies result from the combinatorial composition of the individual types. For example, the type !TC–!TC comprises four distinct, equally probable cases: O–O, O–K, K–O, and K–K, where O and K denote original and knockoff features, respectively. The ratio between original-original interactions and knockoff-involving interactions is therefore $1:3$. In comparison, the type TC–!TC only has two equally probable cases: TC-original and TC-knockoff, which reflects a $1:1$ ratio for this type.
However, the number of false positives cannot be determined directly because #$\#(\text{Knockoff}_{\text{TC--!TC}})$ and #$\#(\text{Knockoff}_{\text{!TC--!TC}})$ are not explicitly measurable. Although knockoff-involving interactions can be identified by the presence of at least one knockoff feature, it is generally not possible to categorize them as either type TC–!TC or !TC–!TC, because original features cannot be definitively classified as correct or incorrect. Therefore, #$\#(\text{Knockoff}_{\text{TC--!TC}})$ and #$\#(\text{Knockoff}_{\text{!TC--!TC}})$ have to be estimated from the subset of knockoff-involving interactions constituted by two knockoff features (i.e., K–K). Specifically, the expected count of !TC–!TC can be estimated from K–K interactions as follows because the random interactions are distributed equally among the knockoff-involving types O–K, K–O, and K–K:
$\widehat{E}\left[\#(\text{Knockoff}_{\text{!TC--!TC}})\right]=3\times\#(\text{Knockoff}_{\text{K--K}})$ (15)
Considering that the total number of knockoff-involving interactions #(Knockoff) is the sum of #$\#(\text{Knockoff}_{\text{TC--!TC}})$ and #$\#(\text{Knockoff}_{\text{!TC--!TC}})$, it further follows:
$\widehat{E}\left[\#(\text{Knockoff}_{\text{TC--!TC}})\right]=\#(\text{Knockoff})-3\times\#(\text{Knockoff}_{\text{K--K}})$ (16)
The expected number of false positives can now be estimated using the knockoff-involving identification calculations from Eq. 15 and Eq. 16: By plugging these estimates into Eq. 14 and using the results to compute the total number of false positives in Eq. 13, we have
$\widehat{E}\left[\#(\text{FP})\right]=\#(\text{Knockoff})-2\times\#(\text{Knockoff}_{\text{K--K}})$ (17)
Appendix A.5 Simulated dataset
Following [59], we use simulation datasets to evaluate Diamond. 10 simulation functions (described in Tab. A.1) are used to generate the simulation datasets.
Table A.1: A test suite of data-generating simulation functions proposed by [59].
@@ Appendix A.5 Simulated dataset
We discovered that the proposed non-additivity distillation procedure is essential; without it, the FDR cannot be controlled by naively using reported interaction importance from existing methods (Fig. A.2). To gain insight into the FDR control failure, we conducted a qualitative comparison assessing interaction importance before and after non-additivity distillation using the simulation function $F_{1}$ across different ML models (Fig. A.3). The primary cause of the FDR control failure lies in the distribution disparity between original-only interactions and those involving knockoffs. This suggests a violation of the knockoff filter’s assumption in controlling FDR. The proposed non-additivity distillation procedure mitigates the disparity by extracting non-additive interaction effects from the reported interaction importance measures, thereby enhancing the utility of knockoff-involving interactions as a negative control for FDR estimation.
Figure A.2: Evaluation of Diamond without non-additivity distillation on simulated datasets. (a) Diamond with the proposed non-additivity distillation procedure effectively controls the FDR across all ML models. (b) Diamond without the proposed non-additivity distillation procedure fails to control the FDR.
@@
Figure A.3: The non-additivity distillation procedure mitigates the disparity between original-only interactions and those involving knockoffs. (a) The reported interaction importance from existing methods reveals a clear disparity in the distribution of interaction importance between original-only interactions and those involving knockoffs, on simulation function $F_{1}$. The distilled non-additive interactions help mitigate distributional disparities. (b) The reported interaction importance from existing methods shows a clear disparity in the cumulative density function between original-only interactions and those involving knockoffs.
Appendix A.6 Real experiments
@@
Figure A.4: Evaluating Diamond on a real Drosophila enhancer dataset with FT-Transformer models. (a) Diamond is applied with FT-Transformer model to identify important non-additive interactions. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top five interaction annotated. The annotated transcription factors are labeled by their UniProt identifiers. The red stars indicate well-characterized interactions in early Drosophila embryos as ground truth. (b) The top interactions reported by the MLP model are qualitatively evaluated from three aspects: the contribution of feature values to the marginal and interaction importance measure, and the contribution of the marginal importance measures to the interaction importance measure.
@@
Figure A.5: Evaluating Diamond on a real Drosophila enhancer dataset with LightGBM models. (a) Diamond is applied with LightGBM model to identify important non-additive interactions. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top five interaction annotated. The annotated transcription factors are labeled by their UniProt identifiers. The red stars indicate well-characterized interactions in early Drosophila embryos as ground truth. (b) The top interactions reported by the MLP model are qualitatively evaluated from three aspects: the contribution of feature values to the marginal and interaction importance measure, and the contribution of the marginal importance measures to the interaction importance measure.
Figure A.6: Evaluating Diamond on a real Drosophila enhancer dataset with XGBoost models. (a) Diamond is applied with XGBoost model to identify important non-additive interactions. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top five interaction annotated. The annotated transcription factors are labeled by their UniProt identifiers. The red stars indicate well-characterized interactions in early Drosophila embryos as ground truth. (b) The top interactions reported by the MLP model are qualitatively evaluated from three aspects: the contribution of feature values to the marginal and interaction importance measure, and the contribution of the marginal importance measures to the interaction importance measure.